# Representing 3D Shapes With 64 Latent Vectors for 3D Diffusion Models

### v1

@@ Abstract
Constructing a compressed latent space through a variational autoencoder (VAE) is the key for efficient 3D diffusion models. This paper introduces COD-VAE, a VAE that encodes 3D shapes into a COmpact set of 1D latent vectors without sacrificing quality. COD-VAE introduces a two-stage autoencoder scheme to improve compression and decoding efficiency. First, our encoder block progressively compresses point clouds into compact latent vectors via intermediate point patches. Second, our triplane-based decoder reconstructs dense triplanes from latent vectors instead of directly decoding neural fields, significantly reducing computational overhead of neural fields decoding. Finally, we propose uncertainty-guided token pruning, which allocates resources adaptively by skipping computations in simpler regions and improves the decoder efficiency. Experimental results demonstrate that COD-VAE achieves 16$\times$ compression compared to the baseline while maintaining quality. This enables 20.8×20.8$\times$20.8 × speedup in generation, highlighting that a large number of latent vectors is not a prerequisite for high-quality reconstruction and generation.
@@ 1 Introduction
Figure 1: (top) Reconstruction IoU (left) and Rendering-FID of the generation results (right) with varying numbers of latent vectors ($M$). Our COD-VAE outperforms VecSet using 8$\times$ fewer latent vectors, achieving 20.8×20.8$\times$20.8 × generation speedup. (bottom) VecSet with $M=64$ struggles to capture details, while our model accurately reconstructs detailed and complex shapes of the objects.
@@
The core component of VecSet is the cross-attention layer, which directly projects an input point cloud into 1D latent vectors and maps them back to continuous neural fields. These direct mappings, operating as learnable downsampling and interpolation, leads to the latent vectors only with a moderate compression ratio. Due to the limited compression ratio, VecSet still requires sufficiently many latent vectors for high-quality reconstruction, which results in excessive computational costs of 3D generative models. In addition, directly mapping latent vectors to neural fields through cross-attention further increases computational demands, creating another computational bottleneck.
In this paper, we introduce COD-VAE, a VAE that encodes 3D shapes into a COmpact set of 1D vectors with an improved compression ratio. COD-VAE replaces the direct mappings between points and latent vectors with a two-stage autoencoder scheme. This scheme enables our model to construct a significantly compressed latent space, thereby accelerating subsequent diffusion models. It achieves high-quality reconstruction with 16×16$\times$16 × fewer latent vectors, as well as efficiency improvement in nerual fields decoding.
@@ 2 Related work
3D generative models.  Previous methods for 3D generation can be categorized into two main approaches. The first approach employs an optimization-based pipeline with 2D diffusion priors to generate 3D objects [34, 25, 23, 38, 42, 36]. These methods incur extensive computational costs due to the optimization process per object. Conversely, the second approach directly trains generative models on 3D datasets, which have shown promising results with significantly faster inference times. These native 3D generative models have widely exlpored, including GAN-based models [1, 40, 15, 3, 12, 32], autoregressive models [41, 30, 27, 45], and normalizing flows [46, 20, 18]. Recent breakthroughs of diffusion models have also inspired a wave of research in the 3D domain [31, 56, 39, 22, 16, 28], which typically adopt the latent diffusion framework [29].
@@
Compact latent space.  The high computational costs of self-attention has led to efforts in the 2D domain for enhancing latent space compression [47, 5]. TiTok [47] proposes to represent images with highly compressed 1D latent tokens, which are obtained using transformers. DC-AE [5] achieves highly compressed 2D latent space to accelerate diffusion models. We aim to extend this paradigm to the 3D domain, by leveraging learnable point patches and triplanes as intermediate representations. Concurrent to our work, other studies [13, 51] introduce variational tokenizers with next scale prediction for 3D autogressive modeling.
@@ 3.2 COD-VAE architecture
The overall architecture of COD-VAE is illustrated in Figure 2. Given an input point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i}^{N}$, our model encodes 1D latent vectors $\mathcal{Z}=\{\mathbf{z}_{i}\}_{i=1}^{M}$ from this point cloud. In the decoding stage, our latent decoder first decompress $\mathcal{Z}$ along the channel dimension, which is then decoded to dense triplane patches. Finally, we compute continuous occupancy values using the reconstructed triplanes and a shallow MLP. Our uncertainty-guided pruning improves the efficiency of the decoder. The following paragraphs describe details of each component.
@@
Latent decoder.  Our decoding process begins with the latent decoder, which decompresses $\mathcal{Z}$ along the channel dimension. This process involves a linear layer that projects $\mathcal{Z}$ into $D$-dimensional features $\mathcal{F}$, and a stack of self-attention layers to process these features. Note that our latent decoder typically incurs negligible computation costs, as the size of our latent vectors set is small. The latent decoder is trained separately from the other components, with explicit objectives focusing on channel decompression. We provide further details of the two-stage training in Sec. 3.3.
Decoder.  After the latent decoding, our decoder reconstructs dense triplane features from the decompressed vectors $\mathcal{F}^{\prime}$ using transformer blocks. While these intermediate triplanes enable effective and efficient neural fields decoding, they also increase computational costs of the decoder, scaling quadratic to the triplane resolutions. To mitigate this, we introduce an uncertainty-guided token pruning that eliminates redundant computations in simple regions.
The architecture of our decoder is presented in Figure 2. We use a learnable token sequence $\mathbf{e}\in\mathbb{R}^{(\frac{R}{f}\times\frac{R}{f})\times C}$ to represent the positions of triplane tokens, where $R$ is the triplane resolutions and $f$ is the patch size. We initialize the triplane tokens by querying from $\mathcal{F}^{\prime}$ using a cross-attention block. Then the auxiliary head predicts uncertainty values of triplane tokens. We retain only the top $25\%$ tokens with the highest uncertainty values and prune the rest. The remaining tokens tokens are added with a shared mask token, concatenated with $\mathcal{F}^{\prime}$, and processed by ViT-style transformer blocks. Finally, we gather the processed triplane tokens and the pruned ones, which are projected by a linear layer to reconstruct full triplane features.
@@ 3.3 VAE training
Two-stage training.  Instead of jointly training all components, we adopt a two-stage training strategy. In the first stage, we train the autoencoder, a model without the KL block and the latent decoder. In the second stage, we freeze the trained components and optimize the KL block and the latent decoder by minimizing the MSE loss between the outputs of the latent decoder, $\mathcal{F}^{\prime}$, and the encoder outputs, $\mathcal{F}$. We also minimize the KL divergence loss, as well as the reconstruction error for stable training in the second stage. This two-stage training guides the latent decoder to focus on the channel decompression of the latent vectors.
@@ 3.4 Diffusion model
We train a latent diffusion model on the compact latent space constructed by our COD-VAE. Following VecSet [50], we adopt transformer-based denoising networks and use the diffusion formulations from EDM [17]. While we do not explicitly aim to improve the generation process itself, the compact latent space provided by COD-VAE significantly enhances the efficiency of our diffusion models.
@@ 4 Experiments
To verify the effectiveness of our method, we first evaluate the reconstruction quality of our COD-VAE. Then, we assess the results of our diffusion models on both class-conditioned and unconditional generation. We provide more details on the experimental setups in the supplements.
@@
Implementation details.  Our model consists of 4 encoder blocks and 12 decoder layers, with $C=512$ and $D=32$. During training, we sample 4,096 query points uniformly from unit cubes and 4,096 points near the objects’ surfaces. We train the autoencoders for 1,600 epochs on ShapeNet and 300 epochs on Objaverse. The second-stage VAE training follows the same number of epochs. To train our models, we use 4 RTX 4090 GPUs for ShapeNet and 4 A6000 GPUs for Objaverse. Training VAEs takes about 3 days for ShapeNet and 6 days for Objaverse, and training diffusion models takes less than 4 days for both ShapeNet and Objaverse. Our code will be made publicly available.
Figure 5: Qualitative reconstruction results. We deliver the reconstruction results of VAEs.
@@ 4.1 Reconstruction results
Results on ShapeNet.  Table 1 reports the reconstruction results on ShapeNet. Our models, both autoencoders and VAEs, outperform all other methods with only using 64 latent vectors. Even with 32 latent vectors, our models achieve reconstruction quality comparble to VecSet ($M=512$), offering a more efficient yet effective option for diffusion models. In contrast, simply reducing the number of latent vectors of VecSet leads to noticeable degradation in both autoencoders and VAEs. This is also evident in the qualitative results presented in Figure 5 (a). Thanks to the two-stage autoencoding, our models with $M=64$ achieve high-quality reconstruction results, whereas VecSet with $M=64$ fails to represent fine details of the objects.
@@ 4.2 Generation results
To further validate our latent representations, we evaluate the generation performance of diffusion models trained on our latent space. We deliver the results of class-conditioned and unconditional generation, which are conducted on ShapeNet and Objaverse, respectively. We follow VecSet and set the number of sampling steps 18 for ShapeNet. For Objaverse, we increase the sampling steps to 50 due to the complexity of this dataset.
@@
Figure 7: Unconditional generation results of our model. We present the results with $M=64$.
@@ 4.3 Ablation study
To verify the effectiveness of the proposed components, we conduct ablation studies on ShapeNet reconstruction. We report the results of autoencoders with $M=32$.
Encoder design.  We first replace our encoder with the encoder of VecSet, which utilizes the cross-attention to directly map point features to latent vectors and self-attention layers to process latent vectors. We compare encoders with various number of cross-attention layers. As shown in Table 5 (a), our encoder outperforms the models with VecSet’s encoders by a large margin. Using more cross-attention layers does not improve performance beyond a certain point. This highlights the importance of intermediate point patches, which enables higher compression ratio by proressive transformations.
Decoder design.  We also replace our decoder with the cross-attention layer as in VecSet, which directly maps latent vectors to query points. We report the ablation results on the decoder design in Table 5 (b). Our decoder improves the overall reconstruction quality as well as efficiency, compared to that of VecSet. In addition, the intermediate triplane representations significantly reduces computation costs of the neural fields decoding, which accompanies more than 20M query points to reconstruct $128^{3}$ grids. Adding more cross-attention layers could offer better reconstruction quality, but at the cost of excessive computations.
@@
CrossAttn ×$\times 12$ 93.4 0.014 96.5
CrossAttn ×$\times 24$ 93.5 0.014 96.6
@@ 5 Discussion
Uncertainty visualization.  To predicted uncertainty values serve as the importance score of the triplane tokens. As presented in Figure 8, our uncertainty head assigns higher scores to more detailed regions, placing more computations on them. This allows our model to avoid heavy computations for simple parts, and place more computational resources on reconstructing details and thin structures.

### v2

@@ Abstract
Constructing a compressed latent space through a variational autoencoder (VAE) is the key for efficient 3D diffusion models. This paper introduces COD-VAE that encodes 3D shapes into a COmpact set of 1D latent vectors without sacrificing quality. COD-VAE introduces a two-stage autoencoder scheme to improve compression and decoding efficiency. First, our encoder block progressively compresses point clouds into compact latent vectors via intermediate point patches. Second, our triplane-based decoder reconstructs dense triplanes from latent vectors instead of directly decoding neural fields, significantly reducing computational overhead of neural fields decoding. Finally, we propose uncertainty-guided token pruning, which allocates resources adaptively by skipping computations in simpler regions and improves the decoder efficiency. Experimental results demonstrate that COD-VAE achieves 16$\times$ compression compared to the baseline while maintaining quality. This enables 20.8×20.8$\times$20.8 × speedup in generation, highlighting that a large number of latent vectors is not a prerequisite for high-quality reconstruction and generation. The code is available at https://github.com/join16/COD-VAE.
@@ 1 Introduction
Figure 1: (top) Reconstruction IoU (left) and Rendering-FID of the generation results (right) with varying numbers of latent vectors ($M$). Our COD-VAE outperforms VecSet using 16$\times$ fewer latent vectors, achieving 20.8×20.8$\times$20.8 × generation speedup. (bottom) VecSet with $M=64$ struggles to capture details, while our model accurately reconstructs detailed and complex shapes of the objects.
@@
The core component of VecSet is the cross-attention layer, which directly projects an input point cloud into 1D latent vectors and maps them back to continuous neural fields. These direct mappings, operating as learnable downsampling and interpolation, leads to the latent vectors only with a moderate compression ratio. Due to the limited compression ratio, VecSet still requires sufficiently many latent vectors for high-quality reconstruction, resulting in excessive computational costs of 3D generative models. Moreover, directly mapping latent vectors to neural fields through cross-attention creates another computational bottleneck.
In this paper, we introduce COD-VAE, a VAE that encodes 3D shapes into a COmpact set of 1D vectors with an improved compression ratio. COD-VAE replaces the direct mappings between points and latent vectors with a two-stage autoencoder scheme. This scheme enables our model to construct a significantly compressed latent space, thereby accelerating subsequent diffusion models. It achieves high-quality reconstruction with 16×16$\times$16 × fewer latent vectors, as well as efficiency improvement in neural fields decoding.
@@ 2 Related work
3D generative models.  Previous methods for 3D generation can be categorized into two main approaches. The first approach employs an optimization-based pipeline with 2D diffusion priors to generate 3D objects [37, 28, 26, 42, 46, 39]. These methods incur extensive computational costs due to the optimization process per object. Conversely, the second approach directly trains generative models on 3D datasets, which have shown promising results with significantly faster inference times. These native 3D generative models have been widely explored, including GAN-based models [1, 44, 17, 4, 14, 35], autoregressive models [45, 33, 30, 50], and normalizing flows [51, 23, 21]. Recent breakthroughs of diffusion models have also inspired a wave of research in the 3D domain [34, 62, 43, 25, 19, 31], which typically adopt the latent diffusion framework [32].
@@
Compact latent space.  The high computational costs of self-attention has led to efforts in the 2D domain for enhancing latent space compression [52, 6]. TiTok [52] proposes to represent images with highly compressed 1D latent tokens, which are obtained using transformers. DC-AE [6] achieves highly compressed 2D latent space to accelerate diffusion models. We aim to extend this paradigm to the 3D domain, by leveraging learnable point patches and triplanes as intermediate representations. Concurrent to our work, other studies [15, 57] propose compact variational tokenizers with next scale prediction for autogressive generation.
@@ 3.2 COD-VAE architecture
The overall architecture of COD-VAE is illustrated in Figure 2. Given an input point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i}^{N}$, our model encodes 1D latent vectors $\mathcal{Z}=\{\mathbf{z}_{i}\}_{i=1}^{M}$ from this point cloud. In the decoding stage, the latent decoder first decompress $\mathcal{Z}$ along the channel dimension, which is then decoded to dense triplane patches. Finally, we compute continuous occupancy values using the reconstructed triplanes and a shallow MLP. The uncertainty-guided pruning further improves the efficiency of the decoder. The following paragraphs describe details of each component.
@@
Latent decoder.  The decoding process begins with the latent decoder, which decompresses $\mathcal{Z}$ along the channel dimension. It consists of a linear layer that projects $\mathcal{Z}$ into $C$-dimensional features, followed by self-attention layers that produce the decompressed features $\mathcal{F}^{\prime}$. This process typically incurs negligible computation costs, as the size of our latent vector set is small. The latent decoder is trained separately from the other components, with explicit training objectives focusing on channel decompression. We provide further details of the two-stage training in Sec. 3.3.
Decoder.  After the latent decoding, the decoder reconstructs dense triplane features from the decompressed vectors $\mathcal{F}^{\prime}$ using transformer blocks. While these intermediate triplanes enable effective and efficient neural fields decoding, they also increase computational costs of the decoder, scaling quadratic to the triplane resolutions. To mitigate this, we introduce an uncertainty-guided token pruning that eliminates redundant computations in simple regions.
[…]
@@ 3.3 VAE training
Two-stage training.  Instead of jointly training all components, we adopt a two-stage training strategy. In the first stage, we train the autoencoder, a model without the KL block and the latent decoder. In the second stage, we freeze the trained components and optimize the KL block and the latent decoder by minimizing the MSE loss between the outputs of the latent decoder, $\mathcal{F}^{\prime}$, and the encoder outputs, $\mathcal{F}$. We also minimize the KL divergence loss, as well as the reconstruction error for stable training. By encouraging the latent decoder to specialize in channel decompression, the two-stage training strategy improves the accuracy of VAEs.
@@ 3.4 Diffusion model
We train a latent diffusion model on the compact latent space constructed by our COD-VAE. Following VecSet [56], we adopt transformer-based denoising networks and use the diffusion formulations from EDM [20]. While we do not explicitly aim to improve the generation process itself, the compact latent space provided by COD-VAE significantly enhances the efficiency of the diffusion models.
@@ 4 Experiments
To verify the effectiveness of our method, we first evaluate the reconstruction quality of COD-VAE. We then assess the performance of diffusion models on both class-conditioned and unconditional generation tasks. Additional details of the experiments and results are provided in the supplements.
@@
Implementation details.  Our model consists of 4 encoder blocks and 12 decoder layers, with $C=512$ and $D=32$. We set the patch size $f=8$ for the decoder in all experiments. During training, we sample 4,096 query points uniformly from unit cubes and 4,096 points near the objects’ surfaces. We train the autoencoders for 1,600 epochs on ShapeNet and 300 epochs on Objaverse. The second-stage VAE training follows the same number of epochs. Our models are trained using 4 RTX 4090 GPUs for ShapeNet and 4 A6000 GPUs for Objaverse. Training VAEs takes about 3 days for ShapeNet and 6 days for Objaverse, and training diffusion models takes less than 4 days for both datasets.
Figure 5: Qualitative reconstruction results. Comparisons of VAEs on ShapeNet (a) and Objaverse (b) are presented.
@@ 4.1 Reconstruction results
Results on ShapeNet.  Table 1 reports the reconstruction results on ShapeNet. Our models, both autoencoders and VAEs, outperform all other methods with only using 64 latent vectors. Even with 32 latent vectors, our models achieve reconstruction quality comparable to VecSet ($M=512$), offering a more efficient yet effective option for diffusion models. In contrast, simply reducing the number of latent vectors of VecSet leads to noticeable degradation in both autoencoders and VAEs. This is also evident in the qualitative results presented in Figure 5 (a). Thanks to the two-stage autoencoding, our models with $M=64$ achieve high-quality reconstruction results, whereas VecSet with $M=64$ fails to represent fine details of the objects.
@@ 4.2 Generation results
To further validate our latent representations, we evaluate the generation performance of diffusion models trained on our compact latent space. We deliver the results of class-conditioned and unconditional generation, which are conducted on ShapeNet and Objaverse, respectively. We follow VecSet and set the number of sampling steps 18 for ShapeNet. For Objaverse, we increase the sampling steps to 50 as it comprises more diverse and complex objects.
@@
Figure 7: Unconditional generation results. Results generated by our model with $M=64$ are presented.
@@ 4.3 Ablation study
To examine the effects of each proposed component, we conduct ablation studies on ShapeNet reconstruction. All ablations are performed using autoencoders with $M=32$.
Encoder design.  We first replace our encoder with the encoder of VecSet, which utilizes the cross-attention to directly map point features to latent vectors and self-attention layers to process latent vectors. We compare against models with encoders having various number of cross-attention layers. As shown in Table 5 (a), our encoder outperforms the models with VecSet’s encoders by a large margin. Using more cross-attention layers does not improve performance beyond a certain point. This highlights the importance of intermediate point patches, which enables higher compression ratio by progressive transformations.
Decoder design.  We also replace our decoder with the cross-attention layer as used in VecSet, which directly maps latent vectors to query points. Ablation results on the decoder design is presented in Table 5 (b). Our decoder improves the overall reconstruction quality as well as efficiency, compared to that of VecSet. In addition, the intermediate triplane representations significantly reduces computation costs of the neural fields decoding, which accompanies more than 20M query points to reconstruct $128^{3}$ grids. Adding more cross-attention layers could offer better reconstruction quality, but at the cost of excessive computations.
@@
CrossAttn ×12$\times$ 12× 12 93.4 0.014 96.5
CrossAttn ×24$\times$ 24× 24 93.5 0.014 96.6
@@ 5 Discussion
Uncertainty visualization.  The predicted uncertainty values serve as the importance score of the triplane tokens. As presented in Figure 8, our uncertainty head assigns higher scores to more detailed regions, placing more computations on them. This allows our model to avoid heavy computations for simple parts, and place more computational resources on reconstructing details and thin structures.